OpenStudy (anonymous):

√ n+5−√n−10√=1, the value of n is

OpenStudy (anonymous):

is that \(\sqrt{n+5}-\sqrt{n+10}=1\)?

OpenStudy (anonymous):

yes thank you!!!

OpenStudy (anonymous):

okay, add \(\sqrt{n+10}\) to both sides:$$\sqrt{n+5}=1+\sqrt{n+10}\\n+5=(1+\sqrt{n+10})^2=1+n+10+2\sqrt{n+10}\\0=6+2\sqrt{n+10}\\\sqrt{n+10}=-3\\n+10=9\\n=-1$$